OpenStudy (anonymous):

Simplify (5x^-1/2)^-2

OpenStudy (australopithecus):

Rules you should acknowledge when simplifying this problem 1. \[x^{-1} = \frac{1}{x}\] Example: \[(2+x)^{-5} = \frac{1}{(2+x)^{5}}\] 2. \[x^2 = x(x)\] Example: \[(x+2)^{2} = (x+2)(x+2) = x(x) + 2x + 2x + 2(2)\]

OpenStudy (australopithecus):

hope this helps if you have any questions please ask

OpenStudy (anonymous):

100x^2